### v1

@@ 1 Introduction
Many of the countries in the WHO South East Asia Region are close to achieving or have achieved complete elimination of malaria. Maldives and Sri Lanka have eradicated malaria, while Thailand, Timor-Leste, Bhutan, and Nepal aim to do so by 2025 [17]. WHO has set the target for a 90% reduction of the global malaria burden and no new indigenous cases by 2030 [16], but many countries are still far from reaching this target due to a combination of various socioeconomic, ecological, and health system-related factors.
@@
India faces many challenges in the planning of malaria control and prevention strategies and timelines due to the malaria epidemiology in India being complicated across diverse demography, topography, and socio-cultural landscapes [11]. The country has put into effect the National Strategic Plan for Malaria Elimination (2017–2022) [12] and the National Framework for Malaria Elimination (NFME) (2016–2030) [11], which outline strategies such as early diagnosis and prompt treatment, vector control, community involvement, and inter-sectoral collaboration. These efforts are directed toward achieving the national goal of eliminating malaria by 2030, with particular emphasis on 27 high-priority districts where malaria transmission is moderate to high.
Effective malaria control and prevention strategies rely heavily on the ability to accurately predict outbreaks and trends in malaria incidence. Traditional surveillance methods often fall short in providing timely and precise forecasts, thereby necessitating the development of advanced predictive models.
@@
Different methods [18, 20] have been proposed for malaria forecasting to address various challenges. These can be categorized into three categories: mathematical methods [9, 19], machine learning methods [19, 1], and statistical methods [19, 7]. Among these, statistical methods are particularly prominent in malaria forecasting due to their intuitive and robust nature. The most prominent statistical models can be broadly classified into three methods: Time Series Regression/Generalized Estimating Model (GLM)[7, 13], Smoothing methods[7], and Auto-Regressive Integrated Moving Average (ARIMA) [7] methods. Despite previous comparisons among these models, their comparative efficacy in terms of the difference between actual and forecasted malaria cases has not been extensively studied.Time series analysis has emerged as a powerful tool for modeling and forecasting infectious diseases, leveraging historical data to predict future trends.
@@ 2.1 Time series regression models with autocorrelated errors
Time series regression models are widely used in various applications to model the relationship between a dependent variable and one or more independent variables over time. One common issue encountered in such models is autocorrelation in the errors, which violates the assumption of independence of residuals in ordinary least squares (OLS) regression. Ignoring autocorrelation can lead to inefficient estimates and misleading inferences [3, 14].
@@
and $u_{t}$ is white noise. In the above conditional mean $\mu(t,\beta)$ can be used to capture the trend as a function of time by using suitable basis functions to capture nonlinear trend and it can also include a vector of other explanatory variables $\boldsymbol{x}_{t}$ which may themselves change over time. For our applications, we have restricted our attention to modeling the trend using orthogonal polynomials of the form $\mu(t,\beta)=\sum_{k=1}^{m}\beta_{k}\psi_{k}(t)$ where $\psi_{k}(\cdot)$’s denote suitable orthogonal basis functions for $k=1,2,\ldots,m$. More generally, the mean function may be expressed as additive models $\mu(t,\beta)=\sum_{k=1}^{m}s(x_{kt},\gamma_{k})$, where $s(x,\gamma)$ are some smooth functions of the components of covariate vector $\boldsymbol{x}_{t}=(x_{1t},\ldots,x_{pt})^{\prime}$ using B-spilnes or other basis functions. The selection of degree $m$ can be accomplished using cross-validation and information criteria. When residuals from a time series model are autocorrelated, it suggests that the model has not fully captured the dynamics of the data. This autocorrelation can be explicitly modeled to improve the accuracy of the time series model. More generally, the mean function may be expressed as additive models $\mu(t,\beta)=\sum_{k=1}^{m}s(x_{kt},\gamma_{k})$, where $s(x,\gamma)$ are some smooth functions of the components of covariate vector $\boldsymbol{x}_{t}=(x_{1t},\ldots,x_{pt})^{\prime}$ using B-splines or other basis functions [15]. However, for the ease of illustration in this paper we consider only polynomial trend without any covariates. But the methodologies described in the paper can be easily extended to more general trend functions (of time) and other potential predictor variables. Next we provide some basic notions of time series modeling which will be used for our subsequent applications.
@@ 2.1.1 Decomposition of Time Series: Trend, Seasonality, and Residuals
The trend component captures the long-term movement in the data. This is often due to underlying factors such as economic growth, technological advancements, or changes in consumer behavior. Trends can be linear, nonlinear, or even exhibit structural breaks over time [4]. The trend component can be modeled using various techniques, depending on its complexity. For simple trends, a linear model with polynomial terms in time can be used:
@@
where $t$ is time and $\beta_{0},\beta_{1},\dots,\beta_{m}$ are the coefficients to be estimated [6]. Alternatively, an orthogonal polynomials can be used which would aid the selection of the appropriate degree $m$ [10].
In more complex scenarios, non-parametric methods such as smoothing splines or locally weighted regression (LOESS) can be employed [5]. These methods do not assume a specific functional form for the trend, allowing for greater flexibility in capturing nonlinear patterns.
@@
Seasonality refers to regular, periodic fluctuations within a time series, typically driven by seasonal factors like the time of year, day of the week, or time of day. Seasonality is characterized by a fixed and known frequency and is assumed to repeat over time [7]. Seasonality can be modeled using dummy variables or Fourier terms, depending on the periodicity and complexity of the seasonal pattern. For instance, a simple seasonal model with monthly data can be written as:
@@
where $T$ is the period of the seasonality and $K$ is the number of harmonics included [7].
@@
The residual component, often referred to as the ”remainder” or ”irregular” component, is what remains after removing the trend and seasonality from the original time series. Residuals are expected to be random noise, ideally following a white noise process. However, in practice, residuals often exhibit autocorrelation, indicating the presence of unexplained structure in the data. Once the trend and seasonality have been modeled, the remaining residuals can be examined for autocorrelation. If autocorrelation is present, an autoregressive model can be fitted to the residuals. For example, if the residuals follow an AR(p) process, the model can be expressed as:
@@
with $\epsilon_{t}$ following an ARIMA process or another suitable autoregressive model [3]. Although Autoregressive Moving Average (ARMA) models are fundamental tools in time series analysis, Hannan and Rissanen (1982) showed that long autoregressive (AR) models can provide consistent estimates for ARMA parameters when the order is large enough and determined using appropriate criteria such as Akaike’s Information Criterion (AIC). So, we are going to restrict our errors process to only AR models.
@@ 2.1.2 Model Evaluation: Adequacy and Accuracy
Residual analysis is the first step in time series model checking. Residuals, defined as the difference between the observed and fitted values, should ideally behave like white noise—exhibiting no autocorrelation, having constant variance, and following a normal distribution [3].
@@
RMSE penalizes large errors more heavily than smaller ones due to the squaring of errors, making it sensitive to outliers. A lower RMSE indicates a better fit and more accurate predictions [7].
@@
MAPE expresses the accuracy as a percentage, making it easy to interpret. However, MAPE can be problematic when actual values $y_{n+t}$ are close to zero, as it can result in very large values or even undefined behavior [2].
@@ 3.1 Data and preprocessing
The state-level malaria case counts are obtained from the Monthly Malaria Information System (MMIS) from the National Centre Vector Borne Diseases Control (NCVBDC). The data contains monthly malaria case counts of various states from January 2020 to December 2023. Based on the annual parasite index (API), Indian states and union territories are categorised into three strata, Category III consisting of the highest-burden states. It is read from CSV file and converted to a Date type and sorted by date to ensure chronological order. Month information is extracted and converted to dummy variables to account for seasonality in the models. A numeric time column is created to represent the date, which is used in the polynomial trend models.
@@ 3.1.1 Trend Estimation, Seasonal Adjustments and Residual
2:$x$ has at least $36$ data points
@@
This model uses log counts with polynomial trend and seasonal dummies. When transforming back from log counts, the variance of residuals is adjusted using the mean of a log-normal distribution, $E[Y_{t}]=\exp\{\mu_{t}+0.5\sigma^{2}\}$ [8] where $\log Y_{t}\sim N(\mu_{t},\sigma^{2})$, with $\mu_{t}=s_{t}+\beta_{1}t+\beta_{2}t^{2}$.
Model 3
@@ 5.3 Application Workflow
The process starts with data upload and preparation. Users provide a CSV file containing malaria case data, including a ”Date” column and state-specific case counts. The application automatically preprocesses this data, converting the ”Date” column to the correct format and sorting the data chronologically. Following this, users can select a state and a modeling method. The application then fits multiple models to the data, utilizing methods such as Polynomial + Seasonal, Log Polynomial + Seasonal, Lag 12, Lag + Polynomial Trend, and ARIMA.
@@ 6 Conclusion
In conclusion, the results suggest that tailored model selection is essential for optimal accuracy. The R Shiny application developed as part of this research serves as a valuable tool for ongoing malaria surveillance and forecasting efforts. While the present models were applied at the state level, their scalability suggests that they can also be adapted for sub-regional (district-level) data, enhancing their utility for more localized public health interventions. Future research could explore incorporating additional climatic and ecological variables and advanced statistical techniques to further refine predictive accuracy and extend the model’s applicability to sub-regional levels.

### v2

@@ 1 Introduction
Many of the countries in the WHO South East Asia Region are close to achieving or have achieved complete elimination of malaria. Maldives and Sri Lanka have eradicated malaria, while Thailand, Timor-Leste, Bhutan, and Nepal aim to do so by 2025 [WHO2022]. WHO has set the target for a 90% reduction of the global malaria burden and no new indigenous cases by 2030 [WHO2015], but many countries are still far from reaching this target due to a combination of various socioeconomic, ecological, and health system-related factors.
@@
India accounts for an estimated 1.7 million malaria cases in 2022 [WHO2023] with transmission intensity and species prevalence varying significantly across states and districts. The country faces many challenges in the planning of malaria control and prevention strategies and timelines due to the malaria epidemiology in India being complicated across diverse demography, topography, and socio-cultural landscapes [IndiaFramework2016]. The country has put into effect the National Strategic Plan for Malaria Elimination (2017–2022) [IndiaPlan2017] and the National Framework for Malaria Elimination (NFME) (2016–2030) [IndiaFramework2016], which outline strategies such as early diagnosis and prompt treatment, vector control, community involvement, and inter-sectoral collaboration. These efforts are directed toward achieving the national goal of eliminating malaria by 2030, with particular emphasis on 27 high-priority districts where malaria transmission is moderate to high.
Despite improvements in India’s malaria surveillance and national reporting, granular district-level data are not uniformly available or updated in real time, and inconsistencies in reporting quality remain a challenge, particularly in remote or under-resourced areas [Dhamnetiya2015, Ghosh2019]. Moreover, publicly accessible datasets often exclude detailed epidemiological variables or lag behind actual transmission timelines, limiting their use for proactive public health interventions. Given these limitations, state-level forecasting serves as a practical intermediate resolution. It offers a balance between spatial specificity and data availability while supporting public health decision-making at the level where most operational vector control and resource allocation decisions are made. Effective malaria control and prevention strategies rely heavily on the ability to accurately predict outbreaks and trends in malaria incidence. Traditional surveillance methods often fall short in providing timely and precise forecasts, thereby necessitating the development of advanced predictive models.
Different methods [Yu2005, Zinszer2015] have been proposed for malaria forecasting to address various challenges. These can be categorized into three categories: mathematical methods [MacDonald1957, Zinszer2012], machine learning methods [Zinszer2012, Anderson1995], and statistical methods [Zinszer2012, Hyndman2018]. Among these, statistical methods are particularly prominent in malaria forecasting due to their intuitive and robust nature. ML methods such as long short-term memory (LSTM) networks have also become popular tools for disease prediction due to their ability to model complex, non-linear patterns. However, for this study, which focuses on time series forecasting with structured, moderately sized datasets, we deliberately chose traditional statistical models. This decision was guided by several key factors. Beyond their proven practical performance and methodological clarity, the interpretability aspect of statistical models was a primary consideration. Unlike many “black-box” ML approaches, the parameters in models like ARIMA correspond directly to interpretable components such as trend, seasonality, and autoregression. This transparency is vital in a public health context, as it allows policymakers to understand the drivers behind a forecast, fostering greater trust and enabling more informed decision-making.
@@
The most prominent statistical models can be broadly classified into three methods: Time Series Regression/Generalized Linear Model (GLM) [Hyndman2018, Modeling2019], Smoothing methods [Hyndman2018], and Auto-Regressive Integrated Moving Average (ARIMA) [Hyndman2018] methods. Despite previous comparisons among these models, their comparative efficacy in terms of the difference between actual and forecasted malaria cases has not been extensively studied.Time series analysis has emerged as a powerful tool for modeling and forecasting infectious diseases, leveraging historical data to predict future trends.
A key study by Singh et al. [Singh2024] conducted a comprehensive time series analysis of national malaria case trends over the past three decades. By applying ARIMA and Holt’s models, the study evaluates the impact of various public health interventions and presents forecasts aligned with India’s 2030 malaria elimination goals. While informative at a national level, it does not directly address the need for tools suited for decentralized forecasting at state or district levels, which our study seeks to provide. At the district level, Chougale et al. [Chougale2025] explored malaria forecasting in Mumbai using both classical time series models and machine learning algorithms such as random forests and gradient boosting. Their study demonstrated that no single model uniformly outperformed others, especially when data quality and seasonal variation differed across wards. This reinforces our emphasis on methodological flexibility and model selection based on context and data availability. Further highlighting the need for region-specific models, a study by Yadav et al. [Yadav2022] the forecasting efficacy of eight different statistical models—including Generalized Linear Models (GLM) and various SARIMA configurations—in two semi-arid districts of Gujarat. Their findings identified a specific SARIMA model as the most appropriate for the region, underscoring that optimal model selection is highly dependent on local data characteristics. Although this work fortifies the long-standing utility of statistical methods for sub-national malaria prediction, the lack of a non-linear trend component and a seasonality component separate from ARIMA and SARIMA estimates are aspects that are addressed in our study.
@@ 2.1 Time series regression models with autocorrelated errors
Time series regression models are widely used in various applications to model the relationship between a dependent variable and one or more independent variables over time. One common issue encountered in such models is autocorrelation in the errors, which violates the assumption of independence of residuals in ordinary least squares (OLS) regression. Ignoring autocorrelation can lead to inefficient estimates and misleading inferences [box1976time, tsay2010analysis].
@@
and $u_{t}$ is white noise. In the above conditional mean $\mu(t,\beta)$ can be used to capture the trend as a function of time by using suitable basis functions to capture nonlinear trend and it can also include a vector of other explanatory variables $\boldsymbol{x}_{t}$ which may themselves change over time. For our applications, we have restricted our attention to modeling the trend using orthogonal polynomials of the form $\mu(t,\beta)=\sum_{k=1}^{m}\beta_{k}\psi_{k}(t)$ where $\psi_{k}(\cdot)$’s denote suitable orthogonal basis functions for $k=1,2,\ldots,m$. More generally, the mean function may be expressed as additive models $\mu(t,\beta)=\sum_{k=1}^{m}s(x_{kt},\gamma_{k})$, where $s(x,\gamma)$ are some smooth functions of the components of covariate vector $\boldsymbol{x}_{t}=(x_{1t},\ldots,x_{pt})^{\prime}$ using B-spilnes or other basis functions. The selection of degree $m$ can be accomplished using cross-validation and information criteria. When residuals from a time series model are autocorrelated, it suggests that the model has not fully captured the dynamics of the data. This autocorrelation can be explicitly modeled to improve the accuracy of the time series model. More generally, the mean function may be expressed as additive models $\mu(t,\beta)=\sum_{k=1}^{m}s(x_{kt},\gamma_{k})$, where $s(x,\gamma)$ are some smooth functions of the components of covariate vector $\boldsymbol{x}_{t}=(x_{1t},\ldots,x_{pt})^{\prime}$ using B-splines or other basis functions [wood2017gam]. However, for the ease of illustration in this paper we consider only polynomial trend without any covariates. But the methodologies described in the paper can be easily extended to more general trend functions (of time) and other potential predictor variables. Next we provide some basic notions of time series modeling which will be used for our subsequent applications.
@@ 2.1.1 Decomposition of Time Series: Trend, Seasonality, and Residuals
The trend component captures the long-term movement in the data. This is often due to underlying factors such as economic growth, technological advancements, or changes in consumer behavior. Trends can be linear, nonlinear, or even exhibit structural breaks over time [cleveland1990stl]. The trend component can be modeled using various techniques, depending on its complexity. For simple trends, a linear model with polynomial terms in time can be used:
@@
where $t$ is time and $\beta_{0},\beta_{1},\dots,\beta_{m}$ are the coefficients to be estimated [draper1998applied]. Alternatively, an orthogonal polynomials can be used which would aid the selection of the appropriate degree $m$ [narula1979orthogonal].
In more complex scenarios, non-parametric methods such as smoothing splines or locally weighted regression (LOESS) can be employed [cleveland1979robust]. These methods do not assume a specific functional form for the trend, allowing for greater flexibility in capturing nonlinear patterns.
@@
Seasonality refers to regular, periodic fluctuations within a time series, typically driven by seasonal factors like the time of year, day of the week, or time of day. Seasonality is characterized by a fixed and known frequency and is assumed to repeat over time [Hyndman2018]. Seasonality can be modeled using dummy variables or Fourier terms, depending on the periodicity and complexity of the seasonal pattern. For instance, a simple seasonal model with monthly data can be written as:
@@
where $T$ is the period of the seasonality and $K$ is the number of harmonics included [Hyndman2018].
@@
The residual component, often referred to as the “remainder” or “irregular” component, is what remains after removing the trend and seasonality from the original time series. Residuals are expected to be random noise, ideally following a white noise process. However, in practice, residuals often exhibit autocorrelation, indicating the presence of unexplained structure in the data. Once the trend and seasonality have been modeled, the remaining residuals can be examined for autocorrelation. If autocorrelation is present, an autoregressive model can be fitted to the residuals. For example, if the residuals follow an AR(p) process, the model can be expressed as:
@@
with $\epsilon_{t}$ following an ARIMA process or another suitable autoregressive model [box1976time]. Although Autoregressive Moving Average (ARMA) models are fundamental tools in time series analysis, Hannan and Rissanen (1982) showed that long autoregressive (AR) models can provide consistent estimates for ARMA parameters when the order is large enough and determined using appropriate criteria such as Akaike’s Information Criterion (AIC). So, we are going to restrict our errors process to only AR models.
@@ 2.1.2 Model Evaluation: Adequacy and Accuracy
Residual analysis is the first step in time series model checking. Residuals, defined as the difference between the observed and fitted values, should ideally behave like white noise—exhibiting no autocorrelation, having constant variance, and following a normal distribution [box1976time].
@@
RMSE penalizes large errors more heavily than smaller ones due to the squaring of errors, making it sensitive to outliers. A lower RMSE indicates a better fit and more accurate predictions [Hyndman2018].
@@
MAPE expresses the accuracy as a percentage, making it easy to interpret. However, MAPE can be problematic when actual values $y_{n+t}$ are close to zero, as it can result in very large values or even undefined behavior [armstrong1985long].
@@ 3.1 Data and preprocessing
The state-level malaria case counts are obtained from the Monthly Malaria Information System (MMIS) from the National Centre Vector Borne Diseases Control (NCVBDC).At the time of the analysis, the most recent validated data available was up to December 2023. The data contains monthly malaria case counts of various states from January 2020 to December 2023. Based on the annual parasite index (API), Indian states and union territories are categorised into three strata, Category III consisting of the highest-burden states. It is read from CSV file and converted to a Date type and sorted by date to ensure chronological order. Month information is extracted and converted to dummy variables to account for seasonality in the models. A numeric time column is created to represent the date, which is used in the polynomial trend models.
@@ 3.1.1 Trend Estimation, Seasonal Adjustments and Residual
2:$x$ has at least 3636 data points
@@
This model uses log counts with polynomial trend and seasonal dummies. When transforming back from log counts, the variance of residuals is adjusted using the mean of a log-normal distribution, $E[Y_{t}]=\exp\{\mu_{t}+0.5\sigma^{2}\}$ [logcorrection] where $\log Y_{t}\sim N(\mu_{t},\sigma^{2})$, with $\mu_{t}=s_{t}+\beta_{1}t+\beta_{2}t^{2}$.
Model 3 (Lag + Trend)
@@ 5.3 Application Workflow
The process starts with data upload and preparation. Users provide a CSV file containing malaria case data, including a “Date” column and state-specific case counts. The application automatically preprocesses this data, converting the “Date” column to the correct format and sorting the data chronologically. Following this, users can select a state and a modeling method. The application then fits multiple models to the data, utilizing methods such as Polynomial + Seasonal, Log Polynomial + Seasonal, Lag 12, Lag + Polynomial Trend, and ARIMA.
@@
While our R Shiny application is designed to accommodate data at various temporal resolutions—such as hourly, weekly, or monthly-scaling the tool to district-level forecasting introduces several practical limitations such as limited data availability data quality and noise, heterogeneity in reporting systems, etc. To mitigate these challenges, the R Shiny tool includes preprocessing options that allow users to aggregate high-resolution data (e.g., hourly counts) into coarser temporal units (e.g., daily or weekly averages). This aggregation can help stabilize variance and improve model fit when high-frequency noise is dominant. However, users should be mindful that reducing temporal granularity may lead to a loss of short-term dynamics, and such trade-offs should be considered based on the intended use case.
@@ 6 Conclusion
The results suggest that tailored model selection is essential for optimal accuracy. The R Shiny application developed as part of this research serves as a valuable tool for ongoing malaria surveillance and forecasting efforts using univariate time series models. Rolling forecast validation was implemented to mitigate the overfitting of models, but continuous validation with incoming data remains essential to ensure sustained model performance. The modeling approach in this study is currently limited to state-level aggregation; extending this approach to finer geographic units such as districts would require addressing challenges like data sparsity, inconsistent reporting, and variability in surveillance infrastructure. Further, future research could explore potentially important external predictors such as climatic variables, intervention coverage, and socioeconomic conditions. Incorporating such covariates into multivariate models could enhance both predictive accuracy and explanatory power.
7 Funding
This research received no dedicated or external funding.